# Study of the decay pattern of $f_{0}(1370)$ as a $\kappa\bar{\kappa}$ molecular state

### v1

@@ Abstract
Assuming that the $f_{0}(1370)$ is a $\kappa\bar{\kappa}$ molecular state, the partial widths of its various decay channels are calculated, including the two-body decay $K\bar{K}$, $\pi\pi$, $\eta\eta$ and the four-body decay $\rho\rho/\sigma\sigma\to 4\pi$ and $K\bar{K}\pi\pi$. The coupling of $g_{f_{0}(1370)\kappa\bar{\kappa}}\approx 13$ GeV estimated from the Weinberg criterion appears to be significantly underestimated. If this coupling is adjusted to $25\sim 40$ GeV, the total width of $f_{0}(1370)$ can be fitted to the measured value $200\sim 500$ MeV. At the center-of-mass energy $\sqrt{s}=1.37$ GeV, the channels that mainly contribute to the total width are $K\bar{K}$, $\pi\pi$ and $4\pi$ ranked as $\Gamma(K\bar{K})>\Gamma(4\pi)\approx\Gamma(\pi\pi)$ with $g_{f_{0}(1370)\kappa\bar{\kappa}}=35$ GeV. Around $1.37$ GeV, the decay widths of the two-body channels $K\bar{K}$, $\pi\pi$ and $\eta\eta$ remain stable with variation in $\sqrt{s}$, whereas the decay widths of the four-body channels $4\pi$ and $K\bar{K}\pi\pi$ increase continuously with $\sqrt{s}$. Most current data are model-dependent and conflicting, such as the $4\pi$ dominant conclusion and the $K\bar{K}$ to $\pi\pi$ ratios. The current data can not rule out the $\kappa\bar{\kappa}$ assignment for $f_{0}(1370)$. Further reliable theoretical and experimental analyses of $f_{0}(1370)$ are required to reveal its nature.
@@ I Introduction
Quantum Chromodynamics (QCD) was established as the theory of strong interaction almost 50 years ago, however, its low-energy regime, due to the invalidation of the perturbation, is still not well understood, particularly the scalar spectrum below $2$ GeV. The strange mass pattern of lightest scalars $f_{0}(500)$, also called $\sigma$, $a_{0}(980)$, $f_{0}(980)$ contradicts the quark model description, which raises some exotic assignment for them, such as molecular state and tetraquark state [1, 2, 3, 4, 5, 6, 7]. It seems that the scalars above $1$ GeV should be assigned to the lowest $0^{++}$ $q\bar{q}$ nonet, which are expected to be heavier than the lowest pseudoscalars and vectors due to its $^{3}P_{0}$ configuration. However, the set of particles possibly filled into this nonet seems to be supernumerary and their mass ordering is also anomalous as shown in the Review of Particle Physics (RPP) [8]. The properties of the $f_{0}(1500)$-$f_{0}(1370)$ are incompatible with them being the singlet-octet system in $q\bar{q}$ nonet. If $f_{0}(1370)$ is considered as dominantly $n\bar{n}$ state then the $s\bar{s}$ state is expected to be $200\sim 300$ MeV heavier and decay strongly into $K\bar{K}$. While the partial widths of $f_{0}(1500)$ decaying into $\pi\pi$ and $4\pi$ are dominant. Its decay ratio to $K\bar{K}$ is only $8.5\pm 1.0\%$ as reported in RPP [8].
@@
In light of this phenomenology, earlier Amsler and Close [9, 10] suggested that the scalars $f_{0}(1370)$, $f_{0}(1500)$ and $f_{0}(1710)$ could originate from a mixing scheme of the scalar glueball with the scalar $q\bar{q}$ state. Subsequently, numerous works have been devoted to the different glueball mixing schemes [11, 12, 13, 14].
@@
On the other hand, some studies on the meson-meson interactions point out that the $f_{0}(1370)$ can be dynamically generated in the isoscalar $\rho\rho$ $S$-wave interaction [15, 16, 17]. However, when including the coupled-channels of pseudoscalars $\pi\pi$ and $K\bar{K}$ in the isoscalar $\rho\rho$ $S$-wave scattering, the Ref. [18] claim that this $\rho\rho$ state is more possibly the $f_{0}(1500)$ based on the pole position and the partial decay widths. Assigning the $f_{0}(1500)$ and $f_{0}(1710)$ as $\rho\rho$ and $K^{*}\bar{K}^{*}$ molecular states, respectively, Ref. [19] has investigated their plausible isospin partners $a_{0}(1450)$ and $a_{0}(1710)$.
@@
It is worth mentioning that the signal of $f_{0}(1370)$ in experiments is subtle and controversial. In 1966, a measurement of the $4\pi$ spectrum in the $\bar{p}n\to 2\pi^{+}3\pi^{-}$ process [20] found a deviation from the phase space distribution, which is explained by a scalar $\rho\rho$ resonance with $m=1.41$ GeV and $\Gamma=90$ MeV. PDG list this data as an observation of $f_{0}(1370)$ [8]. While the width in this analysis is too narrow compared to those in other analyses of $f_{0}(1370)$, which is about $300$ MeV, but it is consistent with the average width of $f_{0}(1500)$. A reanalysis of $\bar{p}n\to 2\pi^{+}3\pi^{-}$ by Gaspero [21] shows that this reaction is dominated by a $4\pi$ resonance with $M=1386$ MeV, $\Gamma=310$ MeV and quantum number $I^{G}J^{PC}=0^{+}0^{++}$, which decays to both $\rho^{0}\rho^{0}$$(29\%)$ and $\sigma\sigma$$(71)\%$. A similar scalar is found in the $p\bar{p}\to\pi^{+}\pi^{-}3\pi^{0}$ by Crystal Barrel collaboration [22], which both decay through $\rho^{+}\rho^{-}$ and $\sigma\sigma$, but the relative strengths for these two decay modes are different. While it might be natural if the comparison of the ratios is made between $\rho^{0}\rho^{0}$-$\sigma\sigma$ and $\rho^{+}\rho^{-}$-$\sigma\sigma$. The data of $\bar{p}d\to\pi^{-}4\pi^{0}p$ demand two scalar states, the $f_{0}(1370)$ and the $f_{0}(1500)$ [23], and both states have masses and widths compatible with previous findings. The analyses on the Crystal Barrel data of $4\pi$ [23, 24] show that $f_{0}(1370)$ couple strongly to $4\pi$. While the analyses of its two-body final states led to contradictory results. Considering $f_{0}(1500)$ as a $\rho\rho$ molecular then it should dominantly decay to $4\pi$. The $f_{0}(1370)$ signal in the $4\pi$ spectrum still has many ambiguities.
@@
The Crystal Barrel data in Ref. [25, 26, 27] manifest a peak in the $\pi\pi$ and $\eta\eta$ $S$-wave at about $1400$ MeV, as well as a further peak at about $1560$ MeV. Its width varies between $200$ and $700$ MeV, depending on theoretical models. A combined fit [28] to the Crystal Barrel data on $p\bar{p}\to 3\pi^{0}$, $\eta\eta\pi^{0}$ and $\eta\pi^{0}\pi^{0}$, which is fully consistent with several other set of data: CERN-Munich data on $\pi^{-}\pi^{+}\to\pi^{-}\pi^{+}$, GAMS data on $\pi^{-}p\to\pi^{0}\pi^{0}n$, and data from ANL and BNL on $\pi^{-}\pi^{+}\to K\bar{K}$, claimed that the $f_{0}(1370)$ is definitely required by the data on $3\pi^{0}$ and $\eta\pi^{0}\pi^{0}$, as well as $\pi\pi\to K\bar{K}$.
@@
Klempt and Zaitsev [29, 30] question the existence of $f_{0}(1370)$. They argue that in the $\eta\eta$ spectra in $p\bar{p}$ annihilation at higher primary energy [31, 32, 33], $f_{0}(1500)$ is always observed clearly but no such “direct” evidence for $f_{0}(1370)$. However, $f_{0}(1370)$ might have weaker decay rate to the $\eta\eta$ channel compared to $f_{0}(1500)$, and as stated in [33] the resonances near the boundaries of the Dalitz plot suffer from the effect of the anticuts. Five primary sets of data requiring $f_{0}(1370)$ are refitted with suitable Breit-Wigner amplitudes by Bugg to further confirm its existence [34].
@@
For the $pp$ central production, an early measurement of $\pi^{+}\pi^{-}$ system by WA102 Collaboration [40] claimed that it was not possible to describe the data above $1$ GeV without the addition of both the $f_{0}(1370)$ and $f_{0}(1500)$ resonances. A couple channel analysis [41] of the centrally produced $K^{+}K^{-}$ [42] and $\pi^{+}\pi^{-}$ [40] final states also obtain the similar results, and in this analysis the branching ratios to $\pi\pi$ and $K\bar{K}$ of $f_{0}(1370)$ have been determined. A relatively recent data from STAR Collaboration [43] on $pp$ central production of $\pi^{+}\pi^{-}$ and $K^{+}K^{-}$ also claim the possible contribution from $f_{0}(1370)$ in addition to the $f_{0}(1500)$. Based on the WA102 measurements in $pp$ central production, the ratio of $f_{0}(1370)\to K\bar{K}$ to $f_{0}(1370)\to\pi\pi$ is determined to be $0.46\pm 0.19$ [44, 41].
@@
For the production in $J/\psi$ radiative decays, the Ref. [45] found convincing evident for the existence of $f_{0}(1370)$ in $J/\psi$ and $\psi(2S)$radiative decay, although only in the $K\bar{K}$ decay channel. The limited statistics in [45] does not allow them to make a partial-wave analysis, leading to their inability to observe the $f_{0}(1370)$ in its $\pi\pi$ decay due to the strong excitation of $f_{2}(1270)$ in its vicinity. While in the higher-statistics measurements on $J/\psi\to\gamma\pi^{0}\pi^{0}$ [46] and $J/\psi\to\gamma K_{S}K_{S}$ [47] the partial wave analysis both show a clear $0^{++}$ peak around $1.4$ GeV. A coupled-channel analysis [48] including these two data obtain the branching ratio $Br(J/\psi\to\gamma f_{0}(1370)\to\gamma\pi\pi)=(3.8\pm 1)\times 10^{-4}$ and $Br(J/\psi\to\gamma f_{0}(1370)\to\gamma K\bar{K})=(1.3\pm 0.4)\times 10^{-4}$. The authors in Ref. [48] claim that the interference between neighboring states plays a decisive role and the change in the sign of the coupling constant in $\pi\pi$ and $K\bar{K}$ decays for $f_{0}(1500)$ with respect to the $f_{0}(1370)$ “background” has first been noticed in Ref. [49].
@@
For the production in $J/\psi$ hadronic decays, BESIII has measured the $\omega K\bar{K}$ [50], $\omega\pi^{+}\pi^{-}$ [51], $\phi K\bar{K}$ and $\phi\pi\pi$ [52] spectra. The leading-order mechanism of $J/\psi$ hadronic decay to $V(PP)$ suggests that the quark flavors within the recoiled meson $V$ must appear in the resonance of the two pseudoscalars. No signals of $f_{0}(1370)$ or $f_{0}(1500)$ are observed in the $\omega K\bar{K}$ and $\omega\pi^{+}\pi^{-}$ spectra. While same as decaying to $\pi\pi$, a clear signal of $f_{0}(1370)$ in the case of recoiling $\phi$. In the $\phi K\bar{K}$ data, there is a shoulder on its upper side that may be fitted by the interference between $f_{0}(1500)$ and $f_{0}(1710)$, but no conspicuous enhancement of $f_{0}(1370)$ is observed. Therefore, a ratio of $Br(J/\psi\to\phi f_{0}(1370)\to\phi K\bar{K})/Br(J/\psi\to\phi f_{0}(1370)\to\phi\pi\pi)=0.08\pm 0.08$ is presented, this value is much lower than those reported by the Particle Data Group. Ref. [52] explain that by the conspicuous signal in $\pi\pi$ but absence of any corresponding peak in $K^{+}K^{-}$.
Although many analyses have claimed the requirement of $f_{0}(1370)$, the majority of analyses conducted to date rely on fitting sums of Breit-Wigner functions, which may yield results that are dependent on the reactions. Especially for the $f_{0}(1370)$, which is a broad resonance closely neighboring the $f_{0}(1500)$.
@@
Motivated by the phenomenology that the signals of $f_{0}(980)$, $f_{0}(1370)$, and $f_{0}(1790)$ appear successively in the $\pi\pi$ spectrum in $J/\psi$ hadronic decay recoiling $\phi$, where the $s\bar{s}$ pair contributes to the $(PP)$ resonance, we argue that the $f_{0}(1370)$ might be a $\kappa\bar{\kappa}$ molecular state, inspired by the $f_{0}(980)$ and $f_{0}(1790)$ being the plausible $K\bar{K}$ and $K^{*}\bar{K}^{*}$ molecules, respectively. In this work, we aim to investigate the decay pattern for the primary decay channels of $f_{0}(1370)$, including $\pi\pi$, $K\bar{K}$, $\eta\eta$, $4\pi$, and $K\bar{K}\pi\pi$, under the scenario of treating $f_{0}(1370)$ as a $\kappa\bar{\kappa}$ molecular state. In Sec. II, we will present the decay mechanism and give the detailed formalism of the calculation. In Sec. III, we will present the numerical results as well as the discussions. A brief summary will be given in Sec. IV.

### v2

@@ Abstract
Under the hypothesis that the $f_{0}(1370)$ is a $\kappa\bar{\kappa}$ molecular state, we calculate the partial widths of its various decay channels, including the two-body decay $K\bar{K}$, $\pi\pi$, $\eta\eta$ and the four-body decay $\rho\rho/\sigma\sigma\to 4\pi$ and $K\bar{K}\pi\pi$. The coupling of $g_{f_{0}(1370)\kappa\bar{\kappa}}\approx 13$ GeV estimated from the Weinberg criterion yields a width of $f_{0}(1370)$ significantly smaller than the experimental data. By adjusting this coupling to $25\sim 40$ GeV, the total width of $f_{0}(1370)$ can be fitted to the measured value $200\sim 500$ MeV. At the center-of-mass energy $\sqrt{s}=1.37$ GeV, the channels that mainly contribute to the total width are $K\bar{K}$, $\pi\pi$ and $4\pi$ ranked as $\Gamma(K\bar{K})>\Gamma(4\pi)\approx\Gamma(\pi\pi)$ with $g_{f_{0}(1370)\kappa\bar{\kappa}}=35$ GeV. Around $1.37$ GeV, the decay widths of the two-body channels $K\bar{K}$, $\pi\pi$ and $\eta\eta$ remain stable with variation in $\sqrt{s}$, whereas the decay widths of the four-body channels $4\pi$ and $K\bar{K}\pi\pi$ increase continuously with $\sqrt{s}$. Most current data are model-dependent and conflicting, particularly regarding the conclusion of $4\pi$ dominance and the ratio of $K\bar{K}$ to $\pi\pi$ decay widths. The current data can not rule out the $\kappa\bar{\kappa}$ assignment for $f_{0}(1370)$. Further reliable theoretical and experimental analyses of $f_{0}(1370)$ are required to reveal its nature.
@@ I Introduction
Quantum Chromodynamics (QCD) was established as the theory of strong interaction almost 50 years ago. However, its low-energy regime is still not well understood due to the breakdown of perturbation theory, particularly the scalar spectrum below $2$ GeV. The strange mass pattern of the lightest scalars $f_{0}(500)$, also called $\sigma$, $a_{0}(980)$, $f_{0}(980)$ contradicts the quark model description, which raises some exotic assignment for them, such as molecular state and tetraquark state [1, 2, 3, 4, 5, 6, 7]. It seems that the scalars above $1$ GeV should be assigned to the lowest $0^{++}$ $q\bar{q}$ nonet, which are expected to be heavier than the lowest pseudoscalars and vectors due to its $^{3}P_{0}$ configuration. However, the set of particles possibly filled into this nonet seems to be supernumerary and their mass ordering is also anomalous as shown in the Review of Particle Physics (RPP) [8]. The properties of the $f_{0}(1500)$-$f_{0}(1370)$ are incompatible with them being the singlet-octet system in $q\bar{q}$ nonet. If $f_{0}(1370)$ is considered as dominantly $n\bar{n}$ state then the $s\bar{s}$ state is expected to be $200\sim 300$ MeV heavier and decay strongly into $K\bar{K}$. While the partial widths of $f_{0}(1500)$ decaying into $\pi\pi$ and $4\pi$ are dominant. Its decay ratio to $K\bar{K}$ is only $8.5\pm 1.0\%$ as reported in the RPP [8].
@@
In light of this phenomenology, earlier Amsler and Close [9, 10] suggested that the scalars $f_{0}(1370)$, $f_{0}(1500)$ and $f_{0}(1710)$ can originate from a mixing scheme of the scalar glueball with the scalar $q\bar{q}$ state. Subsequently, numerous works have been devoted to the different glueball mixing schemes [11, 12, 13, 14].
@@
On the other hand, some studies on the meson-meson interactions suggest that the $f_{0}(1370)$ may be dynamically generated in the isoscalar $\rho\rho$ $S$-wave interaction [15, 16, 17]. However, when the $\pi\pi$ and $K\bar{K}$ coupled-channels are included, Ref. [18] claims that the $\rho\rho$ resonance aligns more closely with the $f_{0}(1500)$, based on its pole position and partial decay widths. Assigning the $f_{0}(1500)$ and $f_{0}(1710)$ as $\rho\rho$ and $K^{*}\bar{K}^{*}$ molecular states, respectively, Ref. [19] has investigated their plausible isospin partners $a_{0}(1450)$ and $a_{0}(1710)$.
@@
It is worth mentioning that the signal of $f_{0}(1370)$ in experiments is subtle and controversial. In 1966, a measurement of the $4\pi$ spectrum in the $\bar{p}n\to 2\pi^{+}3\pi^{-}$ process [20] found a deviation from the phase space distribution, which is explained by a scalar $\rho\rho$ resonance with mass $m=1.41$ GeV and width $\Gamma=90$ MeV. RPP lists this data as an observation of $f_{0}(1370)$ [8]. The width in this analysis is too narrow compared to those in other analyses of $f_{0}(1370)$, which is about $300$ MeV, but it is consistent with the average width of $f_{0}(1500)$. A reanalysis of $\bar{p}n\to 2\pi^{+}3\pi^{-}$ by Gaspero [21] shows that this reaction is dominated by a $4\pi$ resonance with mass $m=1386$ MeV, width $\Gamma=310$ MeV and quantum number $I^{G}J^{PC}=0^{+}0^{++}$, which decays to both $\rho^{0}\rho^{0}$$(29\%)$ and $\sigma\sigma$$(71)\%$. A compatible scalar was observed in $p\bar{p}\to\pi^{+}\pi^{-}3\pi^{0}$ by Crystal Barrel Collaboration [22], which decay through both $\rho^{+}\rho^{-}$ and $\sigma\sigma$ intermediate states with different relative strengths for these two decay modes. The discrepancy is natural because the comparison is made between the ratios of $\rho^{0}\rho^{0}$$\sigma\sigma$ and $\rho^{+}\rho^{-}$$\sigma\sigma$. Later measurements of $\bar{p}d\to\pi^{-}4\pi^{0}p$ require two scalar states, $f_{0}(1370)$ and $f_{0}(1500)$ [23], with mass and width parameters consistent with prior results. While analyses of Crystal Barrel data of $4\pi$ [23, 24] indicate strong coupling of the $f_{0}(1370)$ to the $4\pi$ final state, studies of its two-body decays have produced highly contradictory results. This is further complicated by the $f_{0}(1500)$, a $\rho\rho$ molecule candidate that is expected to dominate the $4\pi$ spectrum. Therefore, the $f_{0}(1370)$ signal in the $4\pi$ spectrum is still ambiguous.
@@
The Crystal Barrel data in Refs. [25, 26, 27] manifest a peak in the $\pi\pi$ and $\eta\eta$ $S$-wave at about $1400$ MeV, as well as a further peak at about $1560$ MeV. Its width varies between $200$ and $700$ MeV, depending on theoretical models. A combined fit [28] to the Crystal Barrel data on $p\bar{p}\to 3\pi^{0}$, $\eta\eta\pi^{0}$ and $\eta\pi^{0}\pi^{0}$, which is fully consistent with several other set of data: CERN-Munich data on $\pi^{-}\pi^{+}\to\pi^{-}\pi^{+}$, GAMS data on $\pi^{-}p\to\pi^{0}\pi^{0}n$, and data from ANL and BNL on $\pi^{-}\pi^{+}\to K\bar{K}$, claimed that the $f_{0}(1370)$ is definitely required by the data on $3\pi^{0}$ and $\eta\pi^{0}\pi^{0}$, as well as $\pi\pi\to K\bar{K}$.
@@
Klempt and Zaitsev [29, 30] question the existence of $f_{0}(1370)$. They argue that in the $\eta\eta$ spectra in $p\bar{p}$ annihilation at higher primary energy [31, 32, 33], $f_{0}(1500)$ is always observed clearly but no such “direct” evidence of the $f_{0}(1370)$. However, $f_{0}(1370)$ might have weaker decay rate to the $\eta\eta$ channel compared to $f_{0}(1500)$, and as stated in [33] the resonances near the boundaries of the Dalitz plot suffer from the effect of the anticuts. Five primary sets of data requiring $f_{0}(1370)$ are refitted with suitable Breit-Wigner amplitudes by Bugg to further confirm its existence [34].
@@
For the $pp$ central production, an early measurement of $\pi^{+}\pi^{-}$ system by WA102 Collaboration [40] claimed that it was not possible to describe the data above $1$ GeV without the addition of both the $f_{0}(1370)$ and $f_{0}(1500)$ resonances. Similar results were obtained from a coupled-channel analysis [41] of centrally produced $K^{+}K^{-}$ [42] and $\pi^{+}\pi^{-}$ [40] final states. A relatively recent data from STAR Collaboration [43] on $pp$ central production of $\pi^{+}\pi^{-}$ and $K^{+}K^{-}$ also claim the possible contribution from $f_{0}(1370)$ in addition to the $f_{0}(1500)$. Based on the WA102 measurements in $pp$ central production, the ratio of $f_{0}(1370)\to K\bar{K}$ to $f_{0}(1370)\to\pi\pi$ is determined to be $0.46\pm 0.19$ [44, 41].
@@
For the production in $J/\psi$ radiative decays, Ref. [45] found convincing evidence of the existence of $f_{0}(1370)$ in $J/\psi$ and $\psi(2S)$ radiative decay, although only in the $K\bar{K}$ decay channel. Due to limited statistics, partial-wave analysis was not performed in [45]. Consequently, the $f_{0}(1370)$ signal in the $\pi\pi$ channel is obscured by the strong excitation of nearby $f_{2}(1270)$. However, partial-wave analyses of higher-statistics $J/\psi\to\gamma\pi^{0}\pi^{0}$ [46] and $J/\psi\to\gamma K_{S}K_{S}$ [47] data both reveal a clear $0^{++}$ structure around $1.4$ GeV. A coupled-channel analysis [48] including these two data obtain the branching ratios $Br(J/\psi\to\gamma f_{0}(1370)\to\gamma\pi\pi)=(3.8\pm 1)\times 10^{-4}$ and $Br(J/\psi\to\gamma f_{0}(1370)\to\gamma K\bar{K})=(1.3\pm 0.4)\times 10^{-4}$. The authors in Ref. [48] claim that the interference between neighboring states plays a decisive role and the change in the sign of the coupling constant in $\pi\pi$ and $K\bar{K}$ decays for $f_{0}(1500)$ with respect to the $f_{0}(1370)$ “background” has first been noticed in Ref. [49].
@@
For the production in $J/\psi$ hadronic decays, BESIII has measured the $\omega K\bar{K}$ [50], $\omega\pi^{+}\pi^{-}$ [51], $\phi K\bar{K}$ and $\phi\pi\pi$ [52] spectra. The leading-order mechanism of $J/\psi$ hadronic decay to $V(PP)$ suggests that the quark flavors within the recoiled meson $V$ must appear in the resonance of the two pseudoscalars. No signals of $f_{0}(1370)$ or $f_{0}(1500)$ are observed in the $\omega K\bar{K}$ and $\omega\pi^{+}\pi^{-}$ spectra. However, same as decaying to $\pi\pi$, a clear signal of $f_{0}(1370)$ is observed in the case of recoiling against $\phi$. In the $\phi K\bar{K}$ data, there is a shoulder on its upper side that may be fitted by the interference between $f_{0}(1500)$ and $f_{0}(1710)$, but no conspicuous enhancement of $f_{0}(1370)$ is observed. Therefore, a ratio of $Br(J/\psi\to\phi f_{0}(1370)\to\phi K\bar{K})/Br(J/\psi\to\phi f_{0}(1370)\to\phi\pi\pi)=0.08\pm 0.08$ is presented, which is much lower than those reported by the Review of Particle Physics. As discussed in Ref. [52], this discrepancy arises from the conspicuous signal in the $\pi\pi$ spectrum and the absence of any corresponding peak in the $K^{+}K^{-}$ spectrum.
While numerous studies have advocated for the necessity of the $f_{0}(1370)$, most analyses to date rely on fits using sums of Breit-Wigner functions. This approach often produces reaction-dependent results, a problem particularly acute for the $f_{0}(1370)$ given its large width and proximity to the $f_{0}(1500)$ resonance.
@@
Motivated by the phenomenology that the signals of $f_{0}(980)$, $f_{0}(1370)$, and $f_{0}(1790)$ appear successively in the $\pi\pi$ spectrum in $J/\psi$ hadronic decay recoiling against $\phi$, where the $s\bar{s}$ pair contributes to the $(PP)$ resonance, we argue that the $f_{0}(1370)$ might be a $\kappa\bar{\kappa}$ molecular state, inspired by the $f_{0}(980)$ and $f_{0}(1790)$ being the plausible $K\bar{K}$ and $K^{*}\bar{K}^{*}$ molecules, respectively. In this work, we aim to investigate the decay pattern for the primary decay channels of $f_{0}(1370)$, including $\pi\pi$, $K\bar{K}$, $\eta\eta$, $4\pi$, and $K\bar{K}\pi\pi$, within the framework where the $f_{0}(1370)$ is treated as a $\kappa\bar{\kappa}$ molecular state. In Sec. II, we present the decay mechanism and the detailed formalism for the calculation. In Sec. III, we will present the numerical results as well as the discussions. A brief summary will be given in Sec. IV.